OpenStudy (anonymous):

Use the formula for computing future value using compound interest to determine the value of an account at the end of 5 years if a principal amount of $2,500 is deposited in an account at an annual interest rate of 6% and the interest is compounded daily. Assume there are 365 days in a year. Do not round until the end of your calculations; then, round to the nearest cent as needed.

OpenStudy (kropot72):

The amount earned on principal P after n payment periods at i interest per payment period is: \[\large A _{n}=P(1+i)^{n}\] Plugging in the values given into the formula we get: \[\large A _{(5\times365)}=2500(1+\frac{0.06}{365})^{(5\times365)}=you\ can\ calculate\]

OpenStudy (anonymous):

im still a little confused, is A(5x365) just 1825=2500(1.00)^1825 ?

OpenStudy (anonymous):

?

OpenStudy (kropot72):

@Apmurillo08 You left out the term 0.06/365. That term must be evaluated and added to 1. Then that sum is raised to the power 1825. The amount after 1825 payment periods is given by: 2500(1 + 0.06/365)^1825.